# Tensor-force effects on nuclear matter in relativistic ab initio theory

### v1

@@
Abstract
Within the relativistic Brueckner-Hartree-Fock theory in the full Dirac space, the tensor-force effects on infinite nuclear matter are elucidated by subtracting the matrix elements of tensor forces from the realistic nucleon-nucleon interaction. The tensor-force effects for the binding energy per particle of symmetric nuclear matter (SNM) as well as the symmetry energy are attractive and are more pronounced around the empirical saturation density, while the tensor forces have little impact on the pure neutron matter. By tuning the tensor-force strength, an infinite (negative) scattering length in the spin-triplet channel is found. This locates the dilute SNM with only the ${}^{3}S_{1}$-${}^{3}D_{1}$ channel interaction at the unitary limit. Its ground-state energy is found proportional to the energy of a free Fermi gas with a scaling factor 0.38, revealing good universal properties. This work paves the way to study the tensor-force effects in neutron stars as well as finite nuclei from realistic nucleon-nucleon interactions and highlights the role of the tensor force on the deviation of the nuclear physics to the unitary limit.
@@ Abstract
As a crucial ingredient of the nucleon-nucleon ($NN$) interaction, the tensor force has an important impact on the structural and dynamical properties of the nuclear many-body systems, particularly the properties of exotic nuclei far away from the stability, which are crucial for understanding the nucleonsynthesis in nuclear astrophysics. Many efforts have been devoted to studying the influence of the tensor force in effective $NN$ interaction in nuclear medium within the configuration interaction shell model Otsuka et al. (2020) and the density functional theories (DFTs) Sagawa and Colò (2014). Due to the difference of modeling the effective $NN$ interaction and the difficulty of determining the tensor-force strengths in finite nuclei, the tensor force effects in nuclear physics are not fully settled.
@@
Starting from the realistic $NN$ interactions that are well calibrated in free space, the microscopic ab initio studies provide a more promising way to study the tensor-force effects, since no free parameters are introduced throughout the solution of the many-body methods. Progress has been made on the role of tensor force in the binding of light nuclei Pudliner et al. (1997) and the symmetry energy of nuclear matter Vidaña et al. (2011). Still, insufficient attention has been paid on the studies of tensor-force effects from microscopic ab initio calculations.
Figure 1: (Color online) The binding energies per particle for (a) SNM and (b) PNM calculated with (solid lines and full symbols) and without (dotted lines and empty symbols) tensor forces. (c) The symmetry energy difference obtained by performing the RBHF calculations with and without tensor force. The results from Ref. Wang et al. (2022a) are also shown for comparison. Details can be found in the text.
@@
The relativistic Brueckner-Hartree-Fock (RBHF) theory plays a key role in addressing the challenge of ab initio methods rooted in a relativistic framework Shen et al. (2019a). The strong correlation in many-body wave function induced by the realistic $NN$ interaction $V$ is absorbed into the effective $G$ matrix, which is obtained as the solution of the Thompson equation in medium Brockmann and Machleidt (1990). In addition to the advantages by solving nuclear many-body systems within a relativistic framework Ring (2012), the main strength of the RBHF theory lies in its inherent ability to account for a class of three-body forces which are important for satisfactorily reproducing the empirical saturation properties of nuclear matter Brown et al. (1987); Brockmann and Machleidt (1990). The RBHF theory has been successfully applied to describe the properties of nuclear matter and finite nuclei Wang et al. (2021); Shen et al. (2017).
@@
In the light of the remarkable reliability of the relativistic ab initio theory, the many-body predictions from RBHF theory have been utilized as pseudo-data to constrain the parameters in both non-relativistic and relativistic DFTs. By applying the RBHF theory in neutron drops, a systematic and specific pattern in the evolution of spin-orbit (SO) splittings is observed Shen et al. (2018); Wang et al. (2019), which is found to have strong tensor-force effects. This provides a novel scheme for the determination of the tensor-force strengths and forms an important guide for the microscopic derivations of nuclear DFTs Shen et al. (2019b); Zhao et al. (2020).
@@
In this work, we aim to directly study the tensor-force effects in nuclear many-body systems from relativistic ab initio theory. As a first step we focus on the infinite nuclear matter, which provides an ideal laboratory accessible to different methods that allows detailed benchmark. The RBHF theory in the full Dirac space is adopted, which considers simultaneously the positive- and negative-energy solutions (NESs) of the Dirac equation, thus avoids the uncertainties suffered from the RBHF calculations without NESs. Details can be found in Refs. Wang et al. (2021, 2022b). To identify the effects of tensor forces, we perform the RBHF calculations, in which the tensor forces in different meson-nucleon couplings are removed from the realistic $NN$ interaction by using the non-relativistic reduced operators Wang et al. (2018)
@@
The factor $\mathcal{F}_{\phi}$ in each meson-nucleon coupling can be found in Table 1 of Ref. Wang et al. (2018). This scheme has been initially applied in analyzing the tensor-force effects on the neutron-proton effective mass splitting in neutron-rich matter Wang et al. (2023). To study the tensor-force effects more precisely, a quenching factor $\lambda$ is introduced in Eq. (1) for each coupling channel $\mathcal{F}_{\phi}$ to control the strength of the tensor force. The factor $\lambda=0$ means no tensor force, while $\lambda=1$ means the tensor forces are considered fully as original.
@@
Figure 1(a) shows the binding energies per particle as a function of the density $\rho$ for symmetric nuclear matter (SNM) calculated within the RBHF theory with and without tensor forces. Clearly the tensor-force effects lead to attraction over a wide range of densities and are more significant at lower densities. The characteristics of attraction can be understood from the Thompson equation, which can be written simply in the operator form $G=V+V(Q/e)V+\cdots$. Here $Q$ and $e$ denote the Pauli operator and the energy denominator, respectively. The tensor force in the realistic interaction $V$ makes no contribution to the binding energy of nuclear matter due to the spin saturation. For the second term $V(Q/e)V$, due to the positive definiteness of quadratic terms of $V$ and the negative energy denominator in the presence of the Pauli operator, the contributions of the tensor force in the second term are always attractive. This implies that the contribution of tensor forces in realistic $NN$ interaction to the nuclear matter binding is reflected through higher-order effects. In addition, among the three parameterizations of the Bonn potential Brockmann and Machleidt (1990), the tensor-force effects of Bonn C are the most pronounced. This is consistent with the well known fact that Bonn C has the strongest tensor force Brockmann and Machleidt (1990), as indicated by the largest $D$-state probability of the deuteron.
Figure 2: (Color online) (a) The dimensionless ratio $r_{t}/a_{t}$ in spin-triplet (left axis) and the binding energy for the deuteron $B_{d}$ (right axis) with variation of the tensor force quenching factor $\lambda$. The upper (lower) horizontal dashed line denotes the locale of the unitary (bound) limit. (b) The ground-state energy per particle for SNM calculated with Bonn A, B, and C, where the scattering length $a_{t}$ in spin-triplet channel is tuned to $-100$ fm. Only contributions from the ${}^{3}S_{1}$-${}^{3}D_{1}$ channel are considered. The results of a free Fermi gas with a scaling factor 0.38 is also displayed for comparison.
@@
In Fig. 1(b), we illustrate the tensor-force effects on the binding energies per particle in pure neutron matter (PNM) where the total isospin $T$ for two-nucleon state must equal to $1$. For the Bonn A potential, the calculations with and without tensor force lead to nearly the same results. This indicates that the total tensor-force effects for Bonn A in isospin triplet with $T=1$ are marginal. Similar conclusions can be drawn for Bonn B and Bonn C potentials. This can be understood by the fact that the PNM is dominated by ${}^{1}S_{0}$ partial wave, while the tensor-force contribution appears in $P$ and higher waves.
@@
The results in SNM and PNM allow us to derive the nuclear symmetry energy $E_{\text{sym}}(\rho)$, which is one of the cutting‐edge topics in nuclear physics and astrophysics. We depict the differences between the symmetry energies obtained with and without tensor forces in Fig. 1(c). The tensor force plays an essential role around the empirical saturation density $\rho_{0}=0.16\pm 0.01\ \text{fm}^{-3}$, beyond which the tensor-force effects become weaker. In comparison to our results, the symmetry energy difference obtained in Ref. Wang et al. (2022a) is very similar, where the tensor-force effects are analyzed by performing the RBHF calculations without the off-diagonal $NN$ matrix elements in the coupled spin triplet channels. By using the Hellmann-Feynman theorem from non-relativistic ab initio method Vidaña et al. (2011), the estimation of the tensor-force contribution to symmetry energy is qualitatively consistent (36 MeV at $\rho=0.187\ \text{fm}^{-3}$), despite the quantitative discrepancy which may result from relativistic effects in $NN$ interaction and many-body method.
@@
It should be noted that, starting from the effective $NN$ interaction within the DFTs, where only the tree-level interactions are considered, tensor force plays no role in nuclear matter due to the spin saturation. This clearly reveals the advantages from ab initio studies that the higher-order effects of the tensor force on nuclear matter can be explored.
@@
The nuclear physics is naturally in proximity to the unitary limit where the two-body scattering length is infinite. This can be recognized by the fact that the scattering lengths in ${}^{1}S_{0}$ (spin singlet) and ${}^{3}S_{1}$ (spin triplet) are appreciably large compared to the effective range of the interaction. The system at the unitary limit exhibits universal properties regardless of the nature of the constituent particles and their mutual interactions. In particular, the ground state energy of any fermionic system at the unitary limit is expected to be proportional to the corresponding ones of a non-interacting Fermi gas with the scaling factor called Bertsch parameter Baker (1999).
Many efforts have been devoted to studying the relation of the dilute PNM properties with the unitary limit. At low densities, the neutron-neutron $(nn)$ interaction is dominated by the ${}^{1}S_{0}$ partial wave with the scattering length $a_{s}^{nn}=-18.9$ fm. Therefore, the dilute PNM is similar to a unitary Fermi gas, which is an idealized system of spin-$1/2$ fermions at the unitary limit. It is found that the Bertsch parameter is around 0.3-0.7 Vidaña (2021).
@@
Different to the PNM, the SNM incorporates additionally the spin triplet ${}^{3}S_{1}$ partial wave, where the scattering length $a_{t}=5.4$ fm implying the existence of a bound state for the neutron-proton $(np)$ system, i.e., the deuteron. It is well known that the tensor force plays an important role in the binding energy of the deuteron and leads to the coupling between ${}^{3}S_{1}$ and ${}^{3}D_{1}$ partial waves. In Fig. 2(a) we present the ratio of the effective range $r_{t}$ and the scattering length $a_{t}$ as a function of the strength of the tensor force. With the continuously decreasing of the quenching factor $\lambda$ of the tensor force, the ratio $r_{t}/a_{t}$ also decreases. Note that the running of $r_{t}/a_{t}$ is dominated by the scattering length $a_{t}$, while $r_{t}$ retains in the range of 1.8-5.1 fm. At a certain $\lambda$, the scattering length $a_{t}$ goes to infinity, locating the $np$ system at the unitary limit, which is denoted by the upper horizontal dashed line. This also illustrates that the strong tensor force in the $np$ interaction makes the $np$ system, to some extent, deviate from the unitary limit.
@@
We also show the influence of the tensor force on the binding energy of deuteron in Fig. 2(a). With the decrease of the factor $\lambda$, the binding energy of the deuteron $B_{d}$ also decreases from 2.2246 MeV to zero until meeting the unitary limit. The rapidest case is found for Bonn C, since this potential has the strongest tensor force. The deuteron is not bound any more after the dimensionless ratio $r_{t}/a_{t}$ crosses the locale of the unitary limit and the bound limit is denoted by the lower horizontal dashed line.
@@
It is interesting to study how the dilute SNM with only the coupled spin triplet ${}^{3}S_{1}$-${}^{3}D_{1}$ partial waves behaves at the unitary limit. By tuning the quenching factor of the tensor force $\lambda=0.4029,0.5624$, and $0.6278$ for Bonn A, B, C, respectively, we obtain the scattering length $a_{t}\approx-100$ fm, which is large enough compared to the effective range $r_{t}$ and the average distance between nucleons in the dilute SNM. From the modified Bonn potentials near the unitary limit, we calculate the ground-state energies per particle for SNM by only considering the $NN$ matrix elements in the ${}^{3}S_{1}$-${}^{3}D_{1}$ channel. To avoid the numerical instabilities at extreme low densities, the Thompson equations are solved only once with initial vanishing single-particle potentials. This should be reasonable since the hypothetically converged single-particle potentials are very close to zero at extreme low densities.
In Fig. 2(b), we present the ground-state energy per particle of SNM within the scheme discussed above. The results from Bonn A, B, and C show very good consistency, not sensitive to the details of the interaction. Furthermore, it is found that the results from our calculations can be well simulated with $\xi(E/A)_{\text{free\ gas}}$, where $(E/A)_{\text{free\ gas}}=(3k^{2}_{F})/(10M)$ is the energy per particle of the free Fermi gas and the $\xi=0.38$ is the aforementioned Bertsch parameter. We also calculate the parameter $\xi$ for dilute PNM and SNM in the ${}^{1}S_{0}$ channel, where the scattering lengths are $a_{s}^{nn}=-18.9$ fm and $a_{s}^{np}=-23.7$ fm. The corresponding results are $\xi=0.60$ and $0.42$, implying the convergence of the Bertsch parameter towards the unitary limit. Besides, the value 0.38 is close to the result 0.372 from quantum Monte Carlo calculations Carlson et al. (2011) for unitary Fermi gas, and 0.376 measured in ultra cold atoms Ku et al. (2012). With increasing density, the differences among the three potentials and the deviations with respect to the $0.38(E/A)_{\text{free\ gas}}$ become evident, which means that the universal properties of SNM are getting deteriorated.
@@
In summary, the tensor-force effects on the equation of state of nuclear matter and the symmetry energy have been studied from the realistic Bonn potentials, by invoking the relativistic Brueckner-Hartree-Fock theory with the tensor force matrix elements subtracted. For the binding energies per particle of symmetric nuclear matter (SNM) and the symmetry energy, the tensor-force effects are attractive and are more pronounced around the empirical saturation density. For pure neutron matter, the tensor-force effects are marginal. The impact of the tensor force on the deviation of the $np$ system to the unitary limit has also been studied. By tuning the tensor-force strength, an infinite (negative) scattering length in the spin triplet can be obtained and the dilute SNM is located at the unitary limit. With only the interaction in ${}^{3}S_{1}$-${}^{3}D_{1}$ channel considered, the ground-state energy of dilute SNM is found proportional to that of a free Fermi gas with a scaling factor 0.38, which reveals very good universal properties. This work paves the way to study the tensor-force effects in neutron stars as well as finite nuclei from realistic nucleon-nucleon interactions and highlights the role of the tensor force on the deviation of the nuclear physics to the unitary limit.
Acknowledgements.

### v2

@@
Tensor-force effects on nuclear matter in relativistic ab initio theory
Sibo Wang Department of Physics and Chongqing Key Laboratory for Strongly Coupled Physics, Chongqing University, Chongqing 401331, China    Hui Tong Helmholtz institute for radiation and nuclear physics, University of Bonn, Bonn D-53115, Germany    Chencan Wang School of Physics and astronomy, Sun Yat-Sen University, Zhuhai 519082, China    Qiang Zhao Center for Exotic Nuclear Studies, Institute for Basic Science, Daejeon 34126, Korea    Peter Ring Department of Physics, Technical University of Munich, Garching D-85747, Germany    Jie Meng mengj@pku.edu.cn State Key Laboratory of Nuclear Physics and Technology, School of Physics, Peking University, Beijing 100871, China
(June 3, 2024)
@@
As a crucial ingredient of the nucleon-nucleon ($NN$) interaction, the tensor force has an important impact on the structural and dynamical properties of the nuclear many-body system, particularly the properties of exotic nuclei far away from the stability, which are crucial for understanding the nucleosynthesis in nuclear astrophysics. Many efforts have been devoted to studying the influence of the tensor force in the effective $NN$ interaction in the nuclear medium within the configuration interaction shell model 2020-Otsuka-RevModPhys.92.015002 and the density functional theories (DFTs) 2014-Sagawa-PPNP . Due to the difference of modeling the effective $NN$ interaction and the difficulty of determining the tensor-force strengths in finite nuclei, the tensor-force effects in nuclear physics still need to be fully settled.
@@
Starting from the realistic $NN$ interactions well calibrated in free space, the microscopic ab initio calculations provide a more promising way to study the tensor-force effects since no free parameters are introduced throughout the solution of the many-body methods. Progress has been made on the role of the tensor force in the binding of light nuclei 1997-Pudliner-PhysRevC.56.1720 and the symmetry energy of nuclear matter 2011-Vidana-PhysRevC.84.062801 . Still, insufficient attention has been paid on the studies of tensor-force effects from microscopic ab initio calculations.
@@
The relativistic Brueckner-Hartree-Fock (RBHF) theory plays a key role in addressing the challenge of ab initio methods rooted in a relativistic framework 2019-Shen-PPNP . The strong correlation in the many-body wave function induced by the realistic $NN$ interaction $V$ is absorbed into the effective $G$ matrix, which is obtained as the solution of the Thompson equation in the medium Brockmann1990_PRC42-1965 . In addition to the advantages by solving nuclear many-body systems within a relativistic framework, the main strength of the RBHF theory lies in its inherent ability to account for a class of three-body forces which are important for satisfactorily reproducing the empirical saturation properties of nuclear matter Brockmann1990_PRC42-1965 ; 2008-LiZH-PhysRevC.77.034316 . The RBHF theory has been successfully applied to describe the properties of nuclear matter and finite nuclei WANG-SB2021_PRC103-054319 ; Shen2017PRC ; 2023-ZouWJ-arXiv .
@@
In light of the remarkable reliability of the relativistic ab initio theory, the many-body predictions from RBHF theory have been utilized as pseudo-data to constrain the parameters in both non-relativistic and relativistic DFTs. By applying the RBHF theory in neutron drops, a systematic and specific pattern in the evolution of spin-orbit splittings is observed 2018-Shen-PLB778.344 , which is found to have strong tensor-force effects. This provides a novel scheme for determining the tensor-force strengths and forms an important guide for the microscopic derivations of nuclear DFTs 2019-Shen-PPNP .
@@
In this work, we aim to directly study the tensor-force effects in nuclear many-body systems from relativistic ab initio theory. As a first step we focus on the infinite nuclear matter, which provides an ideal laboratory accessible to different methods that allows detailed benchmark. The RBHF theory in the full Dirac space is adopted, which considers simultaneously the positive- and negative-energy solutions of the Dirac equation, thus avoiding the uncertainties suffered from RBHF calculations without negative-energy states. Details can be found in Refs. WANG-SB2021_PRC103-054319 ; 2022-Wang-SIBO-PhysRevC.106.L021305 . To identify the effects of tensor forces, we perform the RBHF calculations, in which the tensor forces in different meson-nucleon couplings are removed from the realistic $NN$ interaction by the non-relativistic reduced operators 2018-WangZH-PhysRevC.98.034313
@@
The factor $\mathcal{F}_{\phi}$ in each meson-nucleon coupling can be found in Table 1 of Ref. 2018-WangZH-PhysRevC.98.034313 . To study the tensor-force effects more precisely, a quenching factor $\lambda$ is introduced in Eq. (1) for each coupling channel $\mathcal{F}_{\phi}$ to control the strength of the tensor force. The factor $\lambda=0$ means no tensor force, while $\lambda=1$ means the tensor forces are considered fully.
Figure 1: (Color online) Tensor-force effects on nuclear matter. The binding energies per particle of (a) SNM and (b) PNM calculated with (solid lines and full symbols) and without (dotted lines and empty symbols) tensor forces. (c) The symmetry energy difference obtained by performing the RBHF calculations with and without tensor force. The results from Ref. 2022-CCWang-Chin.Phys.C are also shown for comparison. Details can be found in the text.
@@
Figure 1(a) shows the binding energies per particle as a function of the density $\rho$ for symmetric nuclear matter (SNM) calculated within the RBHF theory with and without tensor forces. The tensor-force effects lead to attraction over a wide range of densities and are more significant at lower densities. The characteristics of attraction can be understood from the Thompson equation, which can be written simply in the operator form $G=V+V(Q/e)V+\cdots$. Here $Q$ and $e$ denote the Pauli operator and the energy denominator, respectively. The tensor force in the first term $V$ does not contribute to the binding energy of nuclear matter due to the spin saturation. For the second term $V(Q/e)V$, due to the positive definiteness of quadratic terms of $V$ and the negative energy denominator in the presence of the Pauli operator, the contributions of the tensor force in the second term are always attractive. This implies that the contribution of tensor forces in realistic $NN$ interaction to the nuclear matter binding is reflected through higher-order effects. In addition, among the three parameterizations of the Bonn potential Brockmann1990_PRC42-1965 , the tensor-force effects of Bonn C are the most pronounced. This is consistent with the well-known fact that Bonn C has the strongest tensor force Brockmann1990_PRC42-1965 , as indicated by the most significant $D$-state probability of the deuteron.
@@
In Fig. 1(b), we illustrate the tensor-force effects on the binding energies per particle in pure neutron matter (PNM) where the total isospin $T$ for two-nucleon state must equal $1$. For the Bonn A potential, the calculations with and without tensor forces lead to nearly the same results. This indicates that the total tensor-force effects for Bonn A in the isospin triplet with $T=1$ are marginal. Similar conclusions can be drawn for Bonn B and Bonn C potentials. This can be understood because the PNM is dominated by ${}^{1}S_{0}$ partial wave, while the tensor-force contribution appears in $P$ and higher waves.
@@
The results in SNM and PNM allow us to derive the nuclear symmetry energy $E_{\text{sym}}(\rho)$, which is one of the cutting‐edge topics in nuclear physics and astrophysics. We depict the differences between the symmetry energies obtained with and without tensor forces in Fig. 1(c). The tensor force is essential around the empirical saturation density $\rho_{0}=0.16\pm 0.01\ \text{fm}^{-3}$, beyond which the tensor-force effects weaken. The differences among Bonn potentials show the uncertainties from realistic $NN$ interactions. Compared to our results, the symmetry energy difference obtained in Ref. 2022-CCWang-Chin.Phys.C is very similar, where the tensor-force effects are analyzed by performing the RBHF calculations without the off-diagonal $NN$ matrix elements in the coupled spin-triplet channels, with the pvCD-Bonn potential 2019-WANGCC-Chin.Phys.C as well as the projection techniques adopted. By using the Hellmann-Feynman theorem and a non-relativistic ab initio method with two- and three-body interactions 2011-Vidana-PhysRevC.84.062801 , the estimation of the tensor-force contribution to the symmetry energy is qualitatively consistent (36 MeV at $\rho=0.187\ \text{fm}^{-3}$), despite the quantitative discrepancy which may result from the $NN$ interactions, the many-body method, and the scheme to extract tensor-force effects.
@@
It should be noted that starting from the effective $NN$ interaction within the DFTs, where only the tree-level interactions are considered, tensor force plays no role in nuclear matter due to the spin saturation. This reveals the advantages from ab initio studies that can explore the higher-order effects of the tensor force on nuclear matter.
Figure 2: (Color online) (a) The dimensionless ratio $r_{t}/a_{t}$ in spin triplet (left axis) and the binding energy of the deuteron $B_{d}$ (right axis) with variation of the tensor force quenching factor $\lambda$. The upper (lower) horizontal dashed line denotes the locale of the unitary (bound) limit. (b) The ground-state energy per particle of SNM calculated with Bonn A, B, and C, where the scattering length $a_{t}$ in spin-triplet channel is tuned to $-100$ fm. Only contributions from the ${}^{3}S_{1}$-${}^{3}D_{1}$ channel are considered. The results of a free Fermi gas with a scaling factor 0.38 is also displayed for comparison.
@@
The dilute fermion system at the unitary limit with infinite two-body scattering length is of general interest. The properties of the system exhibit universalities, regardless of the nature of constituent particles and their mutual interactions. In particular, the ground-state energy is proportional to that of a free Fermi gas, $E/A=\xi(E/A)_{\text{free\ gas}}$, where the scaling factor $\xi$ is called the Bertsch parameter 1999-Baker-PhysRevC.60.054311 . Such studies are of close relevance to various systems in condensed matter physics, nuclear physics, etc.
Nuclear physics is naturally in proximity to the unitary limit. This can be recognized because the scattering lengths in ${}^{1}S_{0}$ (spin singlet) and ${}^{3}S_{1}$ (spin triplet) are appreciably large compared to the effective range of the interaction. Many efforts have been devoted to studying the dilute PNM, which is dominated by the ${}^{1}S_{0}$ partial wave with the scattering length $a_{s}^{nn}=-18.9$ fm. By adjusting the mass of the sigma meson in the Bonn potentials, we tune the scattering length $a_{s}^{nn}$ to $-100$ fm and find the Bertsch parameter is 0.576. This value is close to the results from non-relativistic ab initio calculations $2$ The difference from the experimental measurements on ultra-cold atoms, 0.376(4) 2012-Ku-science.1214987 , might be related to the non-zero effective range of the neutron-neutron ($nn$) interaction 2011-Carlson-PhysRevA.84.061602 .
In comparison to the widely studied two-component (spin-$1/2$) unitary Fermi gas, here we generalize to consider a four-component (spin-$1/2$ and isospin-$1/2$) Fermi gas at zero temperature, since the isospin degree of freedom plays a vital role in nuclear physics. It is still not known whether the four-component unitary Fermi gas shares the same Bertsch parameter with the two-component unitary Fermi gas.
@@
Different from the PNM, the SNM incorporates additionally the spin-triplet ${}^{3}S_{1}$ partial wave, where the scattering length $a_{t}=5.4$ fm implying the existence of a bound state for the neutron-proton $(np)$ system, i.e., the deuteron. It is well known that the tensor force plays a vital role in the deuteron’s binding energy and leads to the coupling between ${}^{3}S_{1}$ and ${}^{3}D_{1}$ partial waves. In Fig. 2(a) we present the ratio of the effective range $r_{t}$ and the scattering length $a_{t}$ as a function of the strength of the tensor force. With the continuously decreasing of the quenching factor $\lambda$ of the tensor force, the ratio $r_{t}/a_{t}$ also decreases. Note that the running of $r_{t}/a_{t}$ is dominated by the scattering length $a_{t}$, while $r_{t}$ retains in the range of 1.8-5.1 fm. At a certain $\lambda$, the scattering length $a_{t}$ goes to infinity, locating the $np$ system at the unitary limit, as denoted by the upper horizontal dashed line. This also illustrates that the strong tensor force in the $np$ interaction makes the $np$ system somewhat deviate from the unitary limit.
@@
We also show the tensor-force effects on deuteron’s binding energy in Fig. 2(a). With the decrease of the factor $\lambda$, the binding energy of the deuteron $B_{d}$ also decreases from 2.2246 MeV to zero until meeting the unitary limit. The most rapid case is found for Bonn C, since this potential has the strongest tensor force. The deuteron is not bound any more after the dimensionless ratio $r_{t}/a_{t}$ crosses the locale of the unitary limit and the bound limit is denoted by the lower horizontal dashed line.
@@
It is interesting to study how the dilute SNM with only the coupled spin-triplet ${}^{3}S_{1}$-${}^{3}D_{1}$ partial waves behave at the unitary limit. By tuning the quenching factor of the tensor force $\lambda=0.4029,0.5624$, and $0.6278$ for Bonn A, B, C, respectively, we obtain the scattering length $a_{t}\approx-100$ fm, whose magnitude is large enough compared to the effective range $r_{t}=2.5$ fm. From the modified Bonn potentials, we calculate the ground-state energies per particle of SNM by only considering the $NN$ matrix elements in the ${}^{3}S_{1}$-${}^{3}D_{1}$ channel. The Thompson equations are solved only once with initial vanishing single-particle potentials to avoid the numerical instabilities at extreme low densities. This should be reasonable since the hypothetically converged single-particle potentials are very close to zero at extremely low densities.
In Fig. 2(b), we present the ground-state energy per particle of SNM within the scheme discussed above. The results from Bonn A, B, and C show very good consistency, not sensitive to the details of the interaction. Furthermore, it is found that the results from our calculations are proportional to that of a free Fermi gas for lower densities. The aforementioned Bertsch parameter is determined as $\xi=0.38$. We also obtain the Bertsch parameter $\xi^{\prime}=0.376$ for dilute SNM in the ${}^{1}S_{0}$ channel, where the scattering length of the $np$ ($nn$, proton-proton) system is tuned to $-100$ fm by adjusting the mass of the sigma meson. The almost same Bertsch parameter for dilute SNM in ${}^{1}S_{0}$ and ${}^{3}S_{1}$-${}^{3}D_{1}$ channels may be related to Wigner’s SU(4) symmetry 1937-Wigner-PhysRev.51.106 , since the scatting lengths in the ${}^{1}S_{0}$ and the ${}^{3}S_{1}$-${}^{3}D_{1}$ are the same ($-100$ fm) after the adjustment.
Particularly, in the ${}^{1}S_{0}$ channel, the Bertsch parameter 0.376 for dilute SNM is smaller than the value 0.576 for dilute PNM, due to the additional $np$ interaction. This indicates that the Bertsch parameter of four-component unitary Fermi gas is different from the two-component case. Our calculations for the four-component unitary Fermi gas with spin and isospin can provide valuable reference to other theoretical studies and ultra-cold atom experiments on four-component fermion systems.
@@
In summary, the tensor-force effects on the equation of state of nuclear matter and the symmetry energy have been studied from the realistic Bonn potentials, by invoking the RBHF theory with the tensor-force matrix elements subtracted. For the binding energies per particle of SNM and the symmetry energy, the tensor-force effects are attractive and are more pronounced around the empirical saturation density. For PNM, the tensor-force effects are marginal. The impact of the tensor force on the deviation of the $np$ system to the unitary limit has also been studied. An infinite (negative) scattering length in the spin triplet can be obtained by tuning the tensor-force strength, and the dilute SNM is located at the unitary limit. With only the interaction in the ${}^{3}S_{1}$-${}^{3}D_{1}$ channel considered, the ground-state energy of dilute SNM is found proportional to that of a free Fermi gas with a scaling factor 0.38, which reveals very good universal properties. By applying the method to analyze the tensor-force effects, this work paves the way to study the tensor-force effects in neutron stars as well as finite nuclei from realistic $NN$ interactions. This work also highlights the role of the tensor force on the deviation of the nuclear physics to the unitary limit and provides valuable reference for studies of the four-component unitary Fermi gas.
Conflict of interest
The authors declare that they have no conflict of interest.
Acknowledgments